OpenStudy (anonymous):

How do I find the set of zeros of f(x)=x^3+4x^2+4x?

OpenStudy (anonymous):

x=0 is one zero The other two you get by solving x^2+4x+4=0 (x+2)^2=0 x=-2 (twice)

OpenStudy (anonymous):

where did you get zero from though?

OpenStudy (anonymous):

f(x) = x (x+2)^2 so if x=0, f(x)=0(x+2)^2=0

OpenStudy (anonymous):

All the addends in the expression are multiplied by x at least once